# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v2

@@ IV.2 Axion domain wall
In this limit, the field profile is changing mainly inside the region of low density, and the field profile has an extent of $1/m_{a}$. This exerts a pressure of $\sim m_{a}f_{a}^{2}/r_{\rm NS}$ on the matter inside, again smaller than the pressure difference due to the changing vacuum energy [^2] In the $\mathbb{Z}_{\mathcal{N}}$ model, the vacuum energy difference is $m_{a}^{2}f_{a}^{2}/\mathcal{N}^{2}$, while the domain wall pressure becomes comparable to this vacuum pressure when $r_{\rm NS}\simeq\frac{\mathcal{N}^{2}}{m_{a}}$, that is $m_{a}\simeq 10^{-8}\,{\rm eV}$ for $\varepsilon\simeq 2\times 10^{-7}$. The density $n_{\rm in}$ for which Eq. (24) is satisfied is still much larger than the critical density for which Eq. (11) saturates, since the equation Eq. (25) is only modified by logarithmic corrections..
@@ V Finite size objects at zero pressure
As was first noted in Refs. [37, 45], an object comprised of nuclear matter at $\theta=\pi$ can be stabilized by the negative pressure $m_{a}^{2}f_{a}^{2}$, rather than gravity. These objects, which we call $\pi$-balls, are held together by the attractive axion potential as long as the baryon chemical potential is too small to allow nucleons to escape (see also  [81, 82]). Given the constraint from white dwarf stability of $\varepsilon>2\times 10^{-7}$, a zero-pressure object balanced by electron degeneracy pressure cannot exist. If a neutron star has a central density large enough to be unstable against axion condensation, it is possible that when such a neutron star is disrupted in a binary neutron star merger, $\pi$-balls may be part of the ejecta. Having an axion field surrounding them, these objects may produce a unique electromagnetic signature. Such questions of production and observational signatures of $\pi$-balls will be explored in the future; in this work, we merely comment on their basic properties.
@@
$M_{\rm min}\simeq 8\times 10^{9}\varepsilon^{3/2}\,{\rm MeV}^{4}/m_{a}^{3}=3.3\times 10^{-21}M_{\odot}\left(\frac{\varepsilon}{10^{-6}}\right)^{21/10}\left(\frac{10^{-8}\,{\rm eV}}{m_{a}}\right)^{3}.$ (36)
@@ VI Observable implications and constraints
In this section, we estimate constraints on the exceptionally light axion parameter space from neutron star observations. We focus on three observables based on modifications to the crust of an axion-condensed neutron star: thermal relaxation times of the crust in accreting systems (Sec. VI.1), cooling rates of isolated neutron stars based on a diminished envelope, first considered in Ref. [40], (Sec. VI.2), and fractional moment inertia of the crust as inferred from pulsar glitches (Sec. VI.3). All of these rely only on the understanding of neutron star matter and dynamics below saturation density. In Sec. VI.4, we summarize the constraints in the context of the $Z_{\mathcal{N}}$ model of exceptionally light axions.
@@ VI.1 Constraint from crust thermal relaxation in x-ray binaries
Neutron stars in transiently accreting x-ray binaries go through periods of outburst and quiescence, reaching a quasi-steady state on astronomical timescales [83]. During outburst, matter is rapidly accreted onto the neutron star surface from a companion and pycnonuclear reactions deep in the crust heat the star. On timescales of tens to hundreds of days (see, e.g., Ref. [84]) the crust cools and comes back into thermal equilibrium with the core. Properties of the crust of neutron stars in these x-ray binaries can be inferred by observing the timescale for the crust to thermally relax following an outburst. The thermal timescale is roughly given by $\tau_{\rm th}\simeq C_{V}(\Delta r)^{2}/\kappa$ where $C_{V}$ is the specific heat of the crust, $\kappa$ is the thermal conductivity, and $\Delta r$ is the thickness of the crust [85]. The calculation of $C_{V}$ and $\kappa$ is not straightforward, depending on pasta phases, superfluidity, magnetic field configuration, and the EOS [86, 87, 88] and is likely to be significantly modified in Scenarios B and C where the crust has more electrons and lacks free neutrons. Without making this detailed calculation, we rely on the quadratic dependence of the thermal relaxation timescale on the thickness of the crust and place the conservative constraint that the total crust thickness should be at least 0.2 km, roughly one-fifth of its value in a normal neutron star.
@@ VI.3 Constraint from pulsar glitches
Pulsars are strongly magnetized, rapidly spinning neutron stars that are observed as a highly regular, periodic source of radiation. Some pulsars, however, have been observed to glitch, a phenomenon in which the pulse period rapidly diminishes on timescales of a day. This is typically understood as resulting from a rapid transfer of angular momentum from the core to the neutron superfluid in the crust of a normal neutron star [91, 92, 93]. After including the effects of coupling the solid and superfluid components of the crust, this provides a lower limit on the fractional moment of inertia in the crust. The Vela pulsar provides the strongest constraint, requiring $I_{c}/I\gtrsim 7\%$ [94]. This is a challenging constraint for many nuclear models to satisfy and it is unclear whether an additional reservoir of angular momentum may be required to explain the glitches of the Vela pulsar [95, 96, 97]. As such, although it is possible that the crust alone cannot account for the extreme glitch behavior of the Vela pulsar, we can conservatively exclude fractional moment of inertia less than 0.1%.
In our Scenarios B and C, there is no neutron superfluid when there is an axion-condensed crust since the crust-core phase transition occurs at a baryon density below the neutron drip. Since free neutrons dominate the crustal moment of inertia, these stars have a fractional moment of inertia that is an order of magnitude too small to explain the Vela pulsar glitches. In Scenario A, $0.068>\varepsilon\gtrsim 0.01$ is also excluded for axion mass $m_{a}\gtrsim 10^{-9}\,\mbox{eV}$ because the crust is mostly in the domain wall which is much thinner than a normal neutron star crust.
Figure 17: Fractional crust moment of inertia fraction of a $1.4\,M_{\odot}$ star as a function of $m_{a}$ and $\varepsilon$ in the RMFT model for three values of $K_{\sigma}$. Note the slightly different vertical scale; for smaller $\varepsilon$, the crust moment of inertia fraction stays nearly constant. In the hatched region, the domain wall approximation fails. In the white band, the entire normal crust is present.
Figure 17 shows the fractional of the moment of inertia of a $1.4\,M_{\odot}$ star as a function of $m_{a}$ and $\varepsilon$. For scenarios B and C, the moment of inertia in the crust is strongly suppressed when there is an axion-condensed crust (cf. Fig. 10). Again, the white band shows where the whole normal crust is present and there is no constraint, the hatched region shows where the domain wall approximation fails, and the solid black line divides the region where the domain wall is at the surface of the star and inside the star. The constraint that the fractional moment of inertia exceeds 0.1% is shown for our three scenarios in Fig. 18. Table 7 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, where, as before, the domain wall does not play an important role and the constraint only depends on $\varepsilon$.
Figure 18: Constraint from observation of pulsar glitches in our three scenarios. We conservatively assume that the crust must constitute $\gtrsim 0.1\%$ of the total neutron star moment of inertia. Parameter space as in Fig. 14.
@@
Table 7: Constraint on $\varepsilon$ from crust fractional moment of inertia for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$.
@@
Min. excluded $\varepsilon$ $7.5\times 10^{-3}$ $1.2\times 10^{-2}$ $2.3\times 10^{-2}$
Max. excluded $\varepsilon$ $0.20$ $0.42$ $0.86$
@@ VI.4 $\mathbb{Z}_{\mathcal{N}}$ model
Table 8: Minimum allowed $\mathcal{N}$ in the $\mathbb{Z}_{\mathcal{N}}$ model for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ when the domain wall connects $\theta=0$ and $\theta=\pi$. In Scenario A, the envelope constraint also permits $\mathcal{N}=3$. $\mathcal{N}$ must be an odd integer for the axion to solve the strong-CP problem in these models.

### v3

@@ IV.2 Axion domain wall
In this limit, the field profile is changing mainly inside the region of low density, and the field profile has an extent of $1/m_{a}$. This exerts a pressure of $\sim m_{a}f_{a}^{2}/r_{\rm NS}$ on the matter inside, again smaller than the pressure difference due to the changing vacuum energy [^2] In the $\mathbb{Z}_{\mathcal{N}}$ model, the vacuum energy difference is $m_{a}^{2}f_{a}^{2}/\mathcal{N}^{2}$, while the domain wall pressure becomes comparable to this vacuum pressure when $r_{\rm NS}\simeq\mathcal{N}^{2}/m_{a}$, that is $m_{a}\simeq 10^{-8}\,{\rm eV}$ for $\varepsilon\simeq 2\times 10^{-7}$. The density $n_{\rm in}$ for which Eq. (24) is satisfied is still much larger than the critical density for which Eq. (11) saturates, since the equation Eq. (25) is only modified by logarithmic corrections..
@@ V Finite size objects at zero pressure
As was first noted in Refs. [37, 45], an object comprised of nuclear matter at $\theta=\pi$ can be stabilized by the negative pressure $m_{a}^{2}f_{a}^{2}$, rather than gravity. These objects, which we call $\pi$-balls, are held together by the attractive axion potential as long as the baryon chemical potential is too small to allow nucleons to escape (see also  [85, 86, 82]). Given the constraint from white dwarf stability of $\varepsilon>2\times 10^{-7}$, a zero-pressure object balanced by electron degeneracy pressure cannot exist. If a neutron star has a central density large enough to be unstable against axion condensation, it is possible that when such a neutron star is disrupted in a binary neutron star merger, $\pi$-balls may be part of the ejecta. Having an axion field surrounding them, these objects may produce a unique electromagnetic signature. Such questions of production and observational signatures of $\pi$-balls will be explored in the future; in this work, we merely comment on their basic properties.
@@
$M_{\rm min}\simeq 8\times 10^{9}\varepsilon^{3/2}\,{\rm MeV}^{4}/m_{a}^{3}=2\times 10^{-7}M_{\odot}\left(\frac{\varepsilon}{10^{-1}}\right)^{21/10}\left(\frac{10^{-9}\,{\rm eV}}{m_{a}}\right)^{3}.$ (36)
@@ VI Observable implications and constraints
In this section, we estimate constraints on the exceptionally light axion parameter space from neutron star observations. We focus on three observables based on modifications to the crust of an axion-condensed neutron star: thermal relaxation times of the crust in accreting systems (Sec. VI.1), cooling rates of isolated neutron stars based on a diminished envelope, first considered in Ref. [40], (Sec. VI.2), and fractional moment inertia of the crust as inferred from pulsar glitches (Sec. VI.3). All of these rely only on the understanding of neutron star matter and dynamics below saturation density. In Sec. VI.4, we summarize the constraints in the context of the $\mathbb{Z}_{\mathcal{N}}$ model of exceptionally light axions.
@@ VI.1 Constraint from crust thermal relaxation in x-ray binaries
Neutron stars in transiently accreting x-ray binaries go through periods of outburst and quiescence, reaching a quasi-steady state on astronomical timescales [87]. During outburst, matter is rapidly accreted onto the neutron star surface from a companion and pycnonuclear reactions deep in the crust heat the star. On timescales of tens to hundreds of days (see, e.g., Ref. [88]) the crust cools and comes back into thermal equilibrium with the core. Properties of the crust of neutron stars in these x-ray binaries can be inferred by observing the timescale for the crust to thermally relax following an outburst. The thermal timescale is roughly given by $\tau_{\rm th}\simeq C_{V}(\Delta r)^{2}/\kappa$ where $C_{V}$ is the specific heat of the crust, $\kappa$ is the thermal conductivity, and $\Delta r$ is the thickness of the crust [89].
The specific heat capacity, $C_{V}$, in the crust of a neutron star is influenced by electrons, ions, and neutrons, depending on the ambient temperature. However, the thermal conductivity, $\kappa$, is primarily determined by electrons across all relevant temperatures [90]. For typical crustal temperatures in the range of $10^{7}$-$10^{8}$ K, as found in accreting neutron stars, the electron contribution provides a lower bound for $C_{V}$, while the neutron contribution—typically suppressed due to superfluidity—sets an upper bound.
In Scenario A, since nuclear interactions remain unchanged, the crust’s composition is also relatively unaltered for pressures much larger than the axion pressure. In Scenarios B and C we expect an increase in electron density above the neutron drip pressure of normal nuclear matter with a maximum increase by a factor of $3$-$4$ near the crust-core boundary in Scenario C. In all scenarios, electron conduction is primarily limited by electron-phonon scattering. To estimate uncertainties in electron scattering rates due to electron-phonon interactions in Scenarios B and C, we modified the phonon spectrum for solids with varying nuclear charge $Z$, mass number $A$, and the corresponding plasma frequency, $\omega_{p}\sim(4\pi e^{2}Z^{2}n_{I}/Am_{n})^{1/2}$ [91]. Axion condensation softens the spectrum of phonons, enhancing electron-phonon scattering and suppressing the contribution of phonons to $C_{V}$ (for a review of thermal transport in neutron star crusts, see Ref. [92]).
At equal nuclear pressure (i.e. not including the negative axion pressure) the lack of dripped neutrons generically enhances $\kappa$ in Scenarios B and C due to a significantly larger electron density when nuclei are treated in mean field theory. In order to outweigh the effects of increased electron density, $Z/A$ would need to decrease by about an order of magnitude. At these densities, $A>100$ and including shell effects is unlikely to produce such a massive change. Nonetheless, lacking an ab initio calculation of neutron-rich nuclei in the axion condensed phase, we place a generous uncertainty on this prediction. At low temperatures, the effect on $C_{V}$ from axion condensation is dominated by a decreased contribution from phonons, which as with $\kappa$ is susceptible to modifications from shell effects. For $T\gtrsim 5\times 10^{7}\,\mathrm{K}$ the increased electron density dominates the change to the heat capacity and shell effects are less important for $C_{V}$. When shell effects are neglected, the increase to $C_{V}$ at high temperature is always smaller than the increase to $\kappa$, leading to a net decrease in $C_{V}/\kappa$.
We do not include any free neutrons in the heat capacity in our comparisons as the neutrons form a superfluid at most densities they appear; including them would only strengthen this claim by introducing an additional reservoir of heat. Furthermore, $C_{V}/\kappa$ decreases as a function of nuclear pressure at these densities, so at equal pressure after including the axion pressure, $C_{V}/\kappa$ would be expected to increase in the axion condensed phase at equal total pressure. At the lowest densities, shell effects in nuclei are expected to dominate and it is challenging to predict the exact properties of these nuclei. However, in the range of $\varepsilon$ that we constrain with thermal relaxation, the part of the crust with positive pressure is all above the normal neutron drip pressure and these shell effects are only dominant in the thin domain wall.
To establish a bound on axion condensation from crust relaxation—without performing detailed self-consistent calculations of $C_{V}$ and $\kappa$—we exploit the quadratic dependence of the thermal relaxation timescale on crust thickness. As a conservative constraint, we require that the total crust thickness be at least 0.2 km, which is approximately one-fifth of its value in a normal neutron star. This assumption effectively limits the increase of $C_{V}/\kappa$ to no more than a factor of 25 when $\theta=\pi$, larger than the maximum increase in $C_{V}/\kappa$ of a factor of 10 at $\theta=\pi$ suggested by the preceding arguments.
@@ VI.3 Constraint from pulsar glitches
Pulsars are spinning neutron stars that emit regular pulses of EM radiation at twice their spin frequency. Pulsar timing studies since the 1970s have cataloged about 3000 sources that exhibit remarkable stability. Precise measurements of the spin period and its steady rate of change have provided useful insights about neutron star populations and their magnetic fields. Some pulsars, however, have been observed to glitch, a phenomenon in which the rotation frequency rapidly increases on timescales of less than a minute. For a concise review of pulsar timing and glitches see [95].
In crust-based models of neutron star glitches, the phenomenon is thought to result from a sudden transfer of angular momentum from an interior superfluid that stores angular momentum in quantized vortices to the crust. As the star gradually spins down, the superfluid cannot adjust smoothly because its vortices are pinned—or otherwise hindered—by a lattice of neutron-rich nuclei [96, 97, 98, 99]. Over time, this creates a buildup of angular momentum stress between the superfluid and the rest of the star. Eventually, this tension is released in a sudden, dramatic, and still poorly understood event that unpins many vortices at once, rapidly spinning up the star.
The spin-up is proportional to the ratio of the moment of inertia of the superfluid component, denoted by $I_{s}$, to the moment of inertia of the star, denoted by $I$. The Vela pulsar places the most stringent constraint on glitch models, requiring that the ratio of superfluid moment of inertia to the total moment of inertia lies between 1.6% and 7% [100]. While there remains ongoing debate about whether nuclear physics models can account for values at the upper end of this range [101, 100, 102]—and speculation about additional angular momentum reservoirs being necessary to fully explain Vela’s glitches [103]—the conventional model, involving a neutron superfluid in the inner crust, remains both natural and viable.
The glitches of the Vela pulsar are challenging to describe in terms of other available models for glitch behavior. Starquakes are a possible alternate explanation for pulsar glitches [104, 105] and recent observations suggest that the Vela pulsar may experience starquakes [106]. However, to explain the high frequency of glitches in the Vela pulsar with starquakes requires either an extreme equation of state in tension with observations of NS masses and radii [107] or results in the Vela pulsar having an unrealistically small mass [108] with the connection between starquakes and glitches still typically requiring the presence of a neutron superfluid [109]. While the possibility of a phase transition deep in the neutron star is potentially relevant for glitches, without any concrete evidence for such a phase transition, this explanation remains speculative.
In the following, we constrain $\varepsilon$ within the framework of the standard glitch model, which requires a substantial superfluid component in the inner crust with $I_{s}/I>1.6\%$. In our Scenarios B and C, when axions condense in the crust, the crust-core phase transition occurs at a baryon density below the neutron drip, and there is no region of coexistence of solid and superfluid matter in the crust. Consequently, glitches cannot be explained in the presence of an axion condensate at the surface of the star when the nuclear force is more attractive. Thus, within the standard glitch models, $\varepsilon<0.27$ is ruled out for Scenario B and $\varepsilon<0.61$ is ruled out for Scenario C provided that $f_{a}$ is not so large that the surface of star has small $\theta$ due to the large domain wall. Even in the domain wall, the increasing nucleon mass brings matter below neutron drip before the interactions become sufficiently repulsive to force neutrons out of nuclei. This result is in part a consequence of the fact that in the ansatz we use in the RMF model, the nuclear force becomes monotonically more attractive as $\theta$ increases. If nuclear matter with moderate isospin asymmetry were only more attractive near $\theta=\pi$ and rapidly became repulsive for intermediate $\theta$, it is possible that enough neutrons could be present in the domain wall to explain glitches when $m_{a}\ll 10^{-8}\,\mathrm{eV}$. This is an issue that warrants further exploration in light of the prediction from our MBPT calculation that neutron matter will be more repulsive at intermediate $\theta$ and the work of past authors [69, 70, 71, 72] that showed that reducing the pion mass made nuclear interactions in neutron matter more repulsive in the vicinity of the physical pion mass (but not at significantly reduced pion mass, as in the axion condensed phase).
We estimate this critical $f_{a}$ by treating the neutron star as a uniform sphere of saturation density matter. The approximation of constant density is conservative since most of the star is at a higher density than the reference density used but is also not particularly impactful since within this approximation $f_{a}^{2}\propto n_{B}^{s}$ and $n_{B}^{s}$ changes by only a factor of at most $3-5$ within the core of a $1.4\,M_{\odot}$ star. Within this constant density approximation, the maximum $f_{a}$ can be found by solving an inhomogeneous Poisson’s equation.
$\nabla^{2}\theta=-\frac{1}{r_{0}^{2}}\frac{\sin\theta}{f(\theta)}$ (38)
where $r_{0}$ is given by
$r_{0}^{-2}=\frac{\sigma_{N}n_{\rm sat}(1+\delta^{\rm(loc)})}{f_{a}^{2}}-m_{a}^{2}\,.$ (39)
In practice, $m_{a}$ has a very small effect on this constraint for $\varepsilon$ that are not very close to the QCD axion line, which this part of the constraint does not touch. In the regime of $m_{a}\ll 1/R_{NS}\simeq 10^{-10}\,\mathrm{eV}$, this can be solved by matching the surface boundary condition to the lowest energy configuration of a massless field, $\theta(r)=\theta_{0}R_{NS}/r$, where $\theta_{0}$ is the critical $\theta$ at which dripped neutrons can appear at positive pressure. The bound on $f_{a}$ we find is below the value necessary for the system to be stable against condensation due to the large extent of the axion field. At the $\varepsilon$ relevant for this part of the constraint, $\beta$-equilibrated matter at or above saturation density is unstable to axion condensation and the system can adiabatically evolve to its ground state. If a more detailed calculation of the nuclear forces shows that $\beta$-equilibrated matter becomes significantly more repulsive for intermediate $\theta$ before becoming attractive at large $\theta$, the exact limits of this bound will need to be revisited but the findings will be qualitatively unchanged.
In Scenario A where dripped neutrons can appear in an axion condensed crust or in any scenario when the critical pressure occurs in the outer crust of a normal star, if the domain wall is large a significant number of neutrons can be present in the domain wall region. For this case, we limit $m_{a}>10^{-8}\,\mathrm{eV}$ so that the domain wall has a thickness of $\mathcal{O}(10\,\mathrm{m})$, more than an order of magnitude smaller than is normally seen. Since the neutron superfluid mechanism for glitches as currently understood requires nearly all of the superfluid neutrons to participate in a normal star, a reduction this severe excludes this possibility. This applies for Scenario A for $0.170.0083<\varepsilon<$ 0.17, for Scenario B for $0.27<0.42$, and for Scenario C for $0.790.61<\varepsilon<$ 0.79.
Lacking a complete and detailed picture of the crust glitch mechanism, we refrain from making a concrete numerical estimate of exactly how many neutrons are necessary to be in tension with observations and rely on the fact that near the boundaries of our constraint the quantity of neutrons changes by many orders of magnitude for relatively small changes in axion parameters. Scenario B and C are particularly notable for the fact that the neutron dripped phase is completely absent for much of the parameter space. The constraint that the quantity of neutrons is reduced by less than an order of magnitude is shown for our three scenarios in Fig. 17. Table 7 shows the constraint on $\varepsilon$ in our three scenarios for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$, where, as before, the domain wall does not play an important role and the constraint only depends on $\varepsilon$. Note that in Fig. 17, there is a small notch in the constraint curves for Scenarios B and C at $m_{a}=10^{-8}\,\mathrm{eV}$. While this is a small effect on the typical log scale for constraint plots, it is a relevant difference when assessing how close constraints come to the QCD axion line at large and small axion masses. It should be noted that only this region of larger axion masses can be confidently constrained within this scheme when using the $\mathbb{Z}_{\mathcal{N}}$ model, which will be discussed in greater detail in the next section.
Figure 17: Constraint from observation of crust-based pulsar glitches in our three scenarios requiring the quantity of neutrons be reduced by less than an order of magnitude. Parameter space as in Fig. 14.
@@
Table 7: Constraint on $\varepsilon$ from observation of crust-based pulsar glitches for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$.
@@
Min. excluded $\varepsilon$ $8.3\times 10^{-3}$ 00 00
Max. excluded $\varepsilon$ $0.17$ $0.42$ $0.79$
@@ VI.4 $\mathbb{Z}_{\mathcal{N}}$ model
A better understanding of the nuclear physics at $\theta=\pi$ is necessary to provide a constraint on the $\mathbb{Z}_{\mathcal{N}}$ model that is independent of our knowledge about how neutron stars are created, and how the axion condensed phase could form following a supernova. On the other hand, it is unclear how a profile with $\theta=\pi$ in the core of a neutron star, and $\theta=(\mathcal{N}-1)\pi/\mathcal{N}$ in most of inner and outer crust could form in the event of neutron star formation in, for example, a core collapse supernova. In an event of a supernova, as the density of the core grows past the critical density in Eq. 11, a region of $\theta=\pi$ emerges in the core, with a domain wall interpolating between the original $\theta=0$ region and the newly formed $\theta=\pi$ region. This domain wall, in particular the part that interpolates between $\theta=(\mathcal{N}-1)\pi/\mathcal{N}$ and $\theta=0$, will need to expand significantly such that most of the inner and outer crust is at $\theta=(\mathcal{N}-1)\pi/\mathcal{N}$ (see Fig. 8 of [37]). During this process, as described in [50], the domain wall gains energy that is $\mathcal{O}(\sigma_{N}/m_{n})$, roughly $5\%$ of the total mass energy of the progenitor star, or at least the total mass energy of the inner and outer crust of the proto-neutron star. This energy, as outlined in [50], will be released slowly, heating up the medium surrounding the neutron star, leading to persistent bright X-ray/optical emissions from the supernova remnant that is not observed. Whereas it is possible that the axion field profile form nearly at rest and are not subsequently accelerated outward, it is hard to imagine how the field profile could gain this much energy through interactions with the baryons, while at the same time, experience exact the right amount of friction to land exactly at the boundary of the outer crust in every single supernova we observed [50, 53]. As a result, we do not consider this scenario further.
@@
Table 8: Minimum allowed $\mathcal{N}$ in the $\mathbb{Z}_{\mathcal{N}}$ model for $m_{a}\gtrsim 10^{-8}\,{\rm eV}$ when the domain wall connects $\theta=0$ and $\theta=\pi$. In Scenario A, the envelope constraint also permits $\mathcal{N}=3$. $\mathcal{N}$ must be an odd integer for the axion to solve the strong-CP problem in these models. Note that these bounds assume the inner region is at $\theta=\pi$, as is preferred in a dynamical formation of the neutron star [50], rather than the critical density minimum at $\theta=(\mathcal{N}-1)\pi/\mathcal{N}$. See text for further discussion.